If Michael Jordan has a vertical leap of 1.89 m, then what is his takeoff speed and his hang time (total time to move upwards to the peak and then return to the ground)? 


1
Expert's answer
2021-12-17T18:14:52-0500

"v=\sqrt{2gh}=6.1~\frac ms,"

"t=2\cdot \sqrt{\frac{2h}g}=1.24~s."